OpenStudy (anonymous):

Find 3 consecutive even numbers where the product of the smaller two numbers is 64 less than the square of the largest number.

OpenStudy (anonymous):

let first number be x second will be x+1 third will be x+2

OpenStudy (anonymous):

now product of smaller numbers is 64 less than square of largest number so you get x(x+1) = (x+2)^2 -64 now solve for x

OpenStudy (anonymous):

I just learned that the variable has to be added x+2 to be evemn

OpenStudy (anonymous):

even*

OpenStudy (anonymous):

there i gave you the equation ^

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

Wrong equation

OpenStudy (blackbird02):

x will be the 1st number x+2 second number x+4 3rd number

OpenStudy (anonymous):

x,x+2,x+4 my bad its x(x+2) = (x+4)^2- 64

OpenStudy (anonymous):

apologies,didnt see the even thing

OpenStudy (blackbird02):

(x)(x+2) = (x+4)^2- 64 simplify them and you have, x^2+2x = x^2 + 8x + 16 - 64

OpenStudy (anonymous):

x = 8

OpenStudy (anonymous):

8,10,12. Thanks guys, I appreciate it.

OpenStudy (blackbird02):

yes, you're welcome